2. Let a and β be multi-indices such that β ≤ a. Prove that
∂^(β) x^(a) = Σ (a,β) β!x^(a - β)
β≤a
1
Expert's answer
2019-10-03T08:52:55-0400

Since it is not indicated that "x\in\mathbb{R}^n" , so I will assume for simplicity that


"\partial^\beta\left(x^\alpha\right)\equiv d^\beta\left(x^\alpha\right)"

Remarks: in the condition there is an expression "\left(\alpha,\,\beta\right)" - this is a binomial coefficient.When solving the problem, I will write it in the traditional form



"\left(\alpha,\,\beta\right)\equiv\dbinom{\alpha}{\beta}\equiv\frac{\alpha!}{\beta!\cdot(\alpha-\beta)!}"

( More information: https://en.wikipedia.org/wiki/Binomial_coefficient )


Next, by definition


"d\left(x^\alpha\right)=\alpha\cdot x^{\alpha-1}\quad\textnormal{Power rule}\\[0.5cm]
d^n\left(f\right)=d^{n-1}\left(d\left(f\right)\right)\equiv d^{n-1}\left(f'\right)"

( More information: https://en.wikipedia.org/wiki/Differentiation_rules )


Then,



"d^\beta\left(x^\alpha\right)=d^{\beta-1}\left(d\left(x^\alpha\right)\right)=
d^{\beta-1}\left(\alpha\cdot x^{\alpha-1}\right)=\\[0.3cm]
=\alpha\cdot d^{\beta-2}\left( d\left(x^{\alpha-1}\right)\right)=
\alpha(\alpha-1)\cdot d^{\beta-2}\left(x^{\alpha-2}\right)=\\[0.3cm]
=\ldots=\alpha(\alpha-1)\cdot\ldots\cdot(\alpha-(\beta-1))\cdot x^{\alpha-\beta}=\\[0.3cm]
=\frac{\alpha(\alpha-1)\cdot\ldots\cdot1}{(\alpha-\beta)\cdot\ldots(1)\cdot\beta!}\cdot \beta!\cdot x^{\alpha-\beta}=\frac{\alpha!}{(\alpha-\beta)!\cdot\beta!}\cdot\beta!\cdot x^{\alpha-\beta}"

Conclusion,



"\boxed{d^\beta\left(x^\alpha\right)=\frac{\alpha!}{(\alpha-\beta)!\cdot\beta!}\cdot\beta!\cdot x^{\alpha-\beta}
\equiv\dbinom{\alpha}{\beta}\cdot\beta!\cdot x^{\alpha-\beta}}"

Q.E.D.